OpenStudy (anonymous):

f(x)=(x-2)^{2}(x-1)^{2}(x+1) - how can i find their zeros and their multiplicity ? i need it too .. thnks for help :)

OpenStudy (anonymous):

You find the zeros when y = 0. Can you draw the equation or try to write it so I can know what exactly it is, I can't read it properly.

OpenStudy (anonymous):

\[f(x)=(x-2)^{2}(x-1)^{2}(x+2)\] this .. thanks

jhonyy9 (jhonyy9):

@bryannjay so how do you think please when will be this function equal zero ?

OpenStudy (anonymous):

After putting it to 0. The equation would look like this: (x−2)^2 (x−1)^2 (x+2) =0. This means that (x-2)^2 = 0, or/and (x-1)^2= 0 or/and (x+2)=0. Then solve it. Good luck.